Six months ago, the exchange rate was 1.20USD per GBP, today it is 1.3USD per GBP. If $1000 were converted to GBP six months ago, how many dollars would there be today is the pounds are converted back today's exchange rate? What has happened to the relative strength of the dollar over the last six months ago. Explain.
Answer #1

GBP purchased 6 months ago = 1000/1.20 = GBP 833.33

Converted into Dollars today = 833.33*1.3 = USD 1083.33

The relative strength of Dollar has decreased i.e. Dollar has weakened against Pound

Since more dollars are now required to buy one Pound